INAF queue observing, 2015 Feb 9-10

Observer: F. Cusano, R. Carini
Telescope Operator: D. Gonzales Huerta
Support Astronomer: Olga Kuhn
Instrument: LUCI/MODS-LBC

Summary

At the sunset the sky is clear.
Some problem with the LUCi rotator at the beginning of the night (solved at UT02:15).
Due to mask exchange problem with LUCI we had to swith to MODs at UT04:00.
The seeing was all night below 1 arcsec. No clouds.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci sdssj0510ori HK , zJ + img in H   COMPLETED
telluric Hip23098 HK and zJ 12 s  
skyflat LUCI H    
Saracco B RDCS-2 red graiting 3 h  
standard Feige 34 red graiting 6 m  
Mignoli MOS22 SDSSJ1148 red graiting 1h 30 m parallel mode
Mignoli LBC_MOS22_red SDSSJ1148 filter Y-FAN 30m  
Mignoli LBC_MOS22_red SDSSJ1148 filter z 30m  
Mignoli LBC_MOS22_red SDSSJ1148 filter i-SLOAN 30m  
Mignoli LBC_MOS22_red SDSSJ1148 filter r-SLOAN 20m  
skyflat LBC filter Y, i   i to be repeated

Details

01:03 The TO opens the dome

01:09 Begin the sky flat field
  • sky flat field in the H filter for Antoniucci program luci.20150210.0011-0031

01.30 problems with the telescope
  • the LUCI's rotator doesn't work, because the fans to cool down the rotator were switched off during the LBTI run.
    TO tries to solve the problem

02:14 problem solved, seeing 0.7
  • We can start the Antoniucci's program sdssj0510ori, start with the imaging
  • ExecuteLUCIscript.sh sdssj0510_H.img luci.20150210.0032-0033 are saturated, the seeing is too low
  • we decrease the exposure time to 4 s, ExecuteLUCIscript.sh sdssj0510_H.img luci.20150210.0034-0039 are saturated
  • the star is very bright, we are defocusing and try again ExecuteLUCIscript.sh sdssj0510_H.img luci.20150210.0040-0045 the images are good

02:29 seeing 0.9, we slew to the spectroscopic acquisition of the target

02:34 try again the acquisition luci.20150210.0046-0049

02:46 ExecuteLUCIscript.sh sdssj0510_ls_HK.spec luci.20150210.0050-0053
  • seeing is 0.6

03:10 ExecuteLUCIscript.sh sdssj0510_ls_zJ.spec luci.0054-57

03:34 We slew to telluric star Hip23098

03:42 ExecuteLUCIscript.sh hip23098_ls_HK.spec luci.0061-0062 S/N low

03:45 we increase the exposure time to increase the S/N

03:47 ExecuteLUCIscript.sh hip23098_ls_zJ.spec luci.0065-0066
  • also in this case the S/N is low, we increase the exposure time

03:50 ExecuteLUCIscript.sh hip23098_ls_zJ.spec luci.0067-0068

03:53 we slew to Piconcelli's A1 program
  • seeing 0.53
  • problem with the mask change

04:23 the problem is not solved by night assistents (also Dave Thompson at the phone), we switch to MODS and slew to the Saracco B program

04:45 Saracco B acqMODS mods.7.RDCSm1.acq mods1r.20150210.0002-0004

04:56 execMODS mods.7.RDCSm1_5.obs mods1r.20150210.0005-0008
  • seeing 0.75

05:46 seeing increase up to 0.9

06:06 execMODS mods.7.RDCSm1_6.obs mods1r.20150210.0009-0012
  • seeing between 0.65 and 0.75

06:43 seeing grows up to 0.85-0.9; we reach the limit of the rotator

07:16 the TO unwraps the rotator
  • seeing 0.9

07:19 acqMODS mods7.recovery.acq
  • mods1r.20150210.0013-0014

07:34 execMODS mods.7.RDCSm1_7.obs
  • mods1r.20150210.0015-0018
  • seeing 0.75

08:16 the sky counts are increasing due to moon.
  • The problem with the LUCI mask graber was solved.

08:45 we slew to standard of Saracco B, feige34
  • acqMODS feige34.acq, mods1r.20150210.0019-0020

08:54 execMODS feige34.obs mods1r.20150210.0021-0023

9:12 slew to Mignoli MOS22
  • acqMODS mods.22.acq mods1r.20150210.0024-0027

09:30 configuration LBC red channel

09:48 run dofpia, /X2, /first, /redonly

09:52 execMODS mods22.4.obs mods1r.20150210.0028-0031
  • the script stop during the offset: 'ERROR:RED go EXPOSURE ALREADY IN PROGRESS'
  • we retry the acquisition but the problem persists
  • problem resolved but we have lost time about 30 m

09:55 run 4_MOS22_Y.ob

10:39 run 5_MOS22_z.ob

10:47 we need to run dofpia

10:56 we run 5_MOS22_z.ob again

11:18 every mods spectra are not good due to initial error in the offset.

11:24 acqMODS mods22.acq mods1r.20150210.0032-0035
  • execMODS mods22.4.obs mods1r.20150210.0036-0039
  • run dofpia

11:29 run 5_MOS22_z.ob again

12:07 run dofpia

12:11 run 6_MOS22_i.ob

12:34 execMODS mods22.5.obs mods1r.20150210.0040-0041

12:50 run dofpia
  • run 7_MOS22_r.ob last files in twilight

13:18 starting sky flat with LBC-R
  • Y filter
  • i filter mostly saturated

14:00 bias LBC

-- %BUBBLESIG{RobertaCarini - 2015-02-10}%
 
Topic revision: r14 - 19 May 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback